OpenStudy (rayford):

Choose the value of the coefficient determinant (D) in the following system. 2x – 5y = –2 x – 3y = 0

OpenStudy (anonymous):

To find the determinant, you first need to place the coefficients in a matrix. \[\left[\begin{matrix}2 & -5 \\ 1 & -3\end{matrix}\right]\] Then, the determinant is given by multiplying along the diagonal down and right, then subtracting the product found by multiplying along the diagonal up and right. So, (2)(-3)-(1)(-5)=-1.

OpenStudy (rayford):

ohhh ok. thank you very much ;]